# Optimal Monotone Depth-Three Circuit Lower Bounds for Majority

### v1

@@ 11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
Note 4.2.
To ensure that each restricted system belongs to ${\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3x,4\}$, we apply induction in the order the properties are stated.
In the next six sections, we present the proof of Theorem 4.1.
@@ 4.2 Proof of the upper bounds for $\Phi_{0}(s,t)$
There exists a pair of elements that appears in at least three distinct $3$-sets.
@@
There exists an element that appears in at least three distinct $3$-sets.
@@
There exists a pair of elements that appears in two distinct $3$-sets.
@@
There exists an element that appears only in one set.
@@
If a $3$-set system $\mathbb{F}$ does not satisfy the conditions $\mathscr{P}(0,2),\mathscr{P}(0,3)$ and $\mathscr{P}(0,4)$, then the following property holds: Let $\{a,b,c\}$ be any set in the system. For $x,y\in\{a,b,c\}$ and $x\neq y$, if there exists sets $S_{x}$, $S_{y}$ where $S_{x}$ contains $x$ and $S_{y}$ contains $y$ and $S_{x}$ and $S_{y}$ are distinct from $\{a,b,c\}$, then $S_{x}$ and $S_{y}$ are disjoint.
@@
If a $3$-set system $\mathbb{F}$ does not satisfy the condition $\mathscr{P}(0,1)$, then every element appears at most twice. If, in addition, the property $\mathscr{P}(0,3)$ does not hold, then by lemma˜4.3, either there exists an $\mathbf{\exists}$} configuration or at least one element appear in exactly one set of $\mathbb{F}$. ∎
Note 4.5.
$\mathscr{P}(0,1)$ is redundant in the sense that every $\mathbb{F}\in{\Xi}_{0}$ satisfies one of the properties, $\mathscr{P}(0,i)$ for $2\leq i\leq 6$. However, we prioritize the application of induction based on $\mathscr{P}(0,1)$ to avoid dealing with restricted systems of three $2$-sets with a common element.
4.2.1 $\mathscr{P}(0,1)$ – There exists a pair of elements that appears in at least three distinct $3$-sets
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,1)$. Let the distinct elements $a$ and $b$ appear together in the sets $\{a,b,c\},\{a,b,d\}$ and $\{a,b,e\}$ where $c$, $d$, and $e$ are distinct elements. Either one of $\{a,b\}$ is such that it always appears in a set together with the other element in $\{a,b\}$ or not. We consider each case in turn and work with the corresponding anchors.
@@ 4.2.1 $\mathscr{P}(0,1)$ – There exists a pair of elements that appears in at least three distinct $3$-sets
Core elements Forced elements Subsystem parameters Inductive bound
@@ 11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
4.2.1.2 $a$ and $b$ each appear in a set without the presence of the other element.
We now extend the anchor by including sets containing either $a$ or $b$ without the other element being present. Observe that when $a=0$ and $b=1$ in the following table, we guaranteed the existence of at least one $2$-set in the resulting restricted system.
Core elements Forced elements Subsystem parameters Inductive bound
@@
4.2.2 $\mathscr{P}(0,2)$ – There exists an element that appears in at least three distinct $3$-sets
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,2)$. Let $a$ be an element that appears in three sets $\{a,b,c\}$, $\{a,d,e\}$ and $\{a,f,g\}$ of $\mathbb{F}$ where $\{b,c\}$, $\{d,e\}$ and $\{f,g\}$ are distinct and do not have a common element.
Core elements Forced elements Subsystem parameters Inductive bound
@@ 4.2.2 $\mathscr{P}(0,2)$ – There exists an element that appears in at least three distinct $3$-sets
4.2.3 $\mathscr{P}(0,3)$ – There exists a pair of elements that appears in two distinct $3$-sets
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,3)$. Assume that $\mathbb{F}$ has a pair $\{a,b\}$ of elements which appears in exactly two sets $\{a,b,c\}$ and $\{a,b,d\}$ since otherwise if $\{a,b\}$ were to appear in three sets, we would have applied induction using $\mathscr{P}(0,1)$. We also know that $a$ (and $b$) cannot appear in any other set since otherwise we would have applied induction using $\mathscr{P}(0,2)$.
Core elements Forced elements Subsystem parameters Inductive bound
@@ 4.2.4 $\mathscr{P}(0,4)$ – There exists a triangle configuration
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,4)$. Let $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ be a triangle configuration in $\mathbb{F}$ where the set $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$. Moreover, we assume that $a$, $b$, and $c$ do not appear in any other set of $\mathbb{F}$.
Core elements Forced elements Subsystem parameters Inductive bound
@@ 4.2.5 $\mathscr{P}(0,5)$ – There exists an $\mathbf{\exists}$} configuration
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,5)$. Let $\{a,b,c\}$ be the spine of the $\mathbf{\exists}$} configuration, and $\{a,d,e\},\{b,f,g\}$ and $\{c,h,i\}$ are its arms where the set $\{d,e,f,g,h,i\}$ is disjoint from $\{a,b,c\}$. We assume $a,b$, and $c$ do not appear in any other sets.
Core elements Forced elements Subsystem parameters Inductive bound
@@
4.2.6 $\mathscr{P}(0,6)$ – There exists an element that appears only in one set
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,6)$. Let $a$ be such that $a$ appears exactly once in $\mathbb{F}$ in the set $\{a,b,c\}$. We consider two cases: Each of $b$ and $c$ appear in a different set or one of them does not appear in any other set.
4.2.6.1 Each of $b$ and $c$ appear in a different set:
@@
Core elements Forced elements Subsystem parameters Inductive bound
@@
4.2.6.2 One of $\{b,c\}$ does not appear in any other set:
Assume that $\mathbb{F}$ contains $\{a,b,c\}$ and $a$ and $b$ do not appear anywhere else. Either $c$ does not appear in any other set or there is a set of the form $\{c,d,e\}$.
Core elements Forced elements Subsystem parameters Inductive bound
@@ 11 $a=b=0$ n3n-3 t1t-1 $s$ 0 $\frac{1}{3}$ $\frac{1}{3}$
When $c=0$, the number of transversals when $c$ appears in another set is dominated by the number of transversals when $c$ does not appear in any other set.
@@ 4.3 Proof of the Upper bounds for $\Phi_{1}(s,t)$
Let $\Xi$} the family of $3$-set systems that contain a $2$-set. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n$ elements such that $\mathbb{F}=\{a,b\}\cup\mathbb{F}^{\prime}$ with threshold $t$ and deficit $s$. The following properties determine the subclasses.
@@
At least one of the elements in $\{a,b\}$ appears at least twice.
4.3.1 At least one variable in $\{a,b\}$ appears exactly once  – $\mathscr{P}(1,1)$
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 0 n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
0 11 n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
@@ 4.3.1 At least one variable in $\{a,b\}$ appears exactly once  – $\mathscr{P}(1,1)$
4.3.2 At least one of the elements in $\{a,b\}$ appears at least twice  – $(\neg\mathscr{P}(1,1))\land\mathscr{P}(1,2)$
In this case, $\mathbb{F}$ has a set of the form $\{a,c,d\}$ where $\{c,d\}$ is disjoint from $\{a,b\}$.
$a$ $b$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
@@ 4.4 Proof of the Upper bounds for $\Phi_{2o}(s,t)$
We now prove the upper bounds for $3$-set systems which contain two overlapping $2$-sets as stated in Theorem 4.1. Let $\Xi$} be the class of $3$-set system with two overlapping $2$-sets. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 3$ elements with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains the overlapping $2$-sets $\{a,b\}$ and $\{a,c\}$.
We first deal with the simpler case of odd $s$. We will show that the number of $t$-transversals is at most $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Odd $s$
11 n1n-1 t1t-1 s2s-2 $\frac{7}{10}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 $\frac{3}{10}$
@@
For the rest of the section, let $s$ be even. We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{2}{3})(\frac{2}{3})^{\frac{s}{2}}3^{t}$. For this purpose, we consider the following subclasses of $\Xi$}.
@@
There are two sets containing both $b$ and $c$.
@@
There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
@@
There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
@@
For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
@@
One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element.
@@
There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and exactly one set that contains both $b$ and $c$.
@@
There are exactly two sets that contain $b$ and exactly two sets that contains $c$ and $b$ and $c$ do not appear together.
Lemma 4.6.
If $\mathbb{F}$ satisfies $\neg(\bigwedge_{j=1}^{3}\mathscr{P}(2o,j))$, then $a$ appears exactly twice and $b$ and $c$ appear at least two times and at most three times in $\mathbb{F}$. Moreover, the following scenarios cover all the situations if one of $b$ or $c$ were to appear three times.
  1. 1.
    There are two sets containing both $b$ and $c$.
  2. 2.
    There is exactly one set in which both $b$ and $c$ appear together and exactly two other sets, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
  3. 3.
    There exist sets $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another set that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
  4. 4.
    For each $x\in\{b,c\}$, there are exactly two sets other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
  5. 5.
    One of $b$ or $c$ appears exactly in three sets which do not contain the other element and the other element appears exactly in two sets which do not contain the first element.
Lemma 4.7.
If $\mathbb{F}$ satisfies $\neg\bigwedge_{j=1}^{8}\mathscr{P}(2o,j)$, then $a$, $b$ and $c$ appear exactly twice.

### v2

@@ 11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
Note 4.3.
To ensure that each restricted CNF belongs to ${\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3,4\}$, we apply induction in the order the properties are stated.
Note 4.4.
In the inductive step, we occasionally have situations in which the exact number of variables in the restricted formula is not guaranteed. In such cases, we inductively use $\Phi_{i}(s,t)\leq\Phi_{i}(s^{\prime},t)$ for every type $i$, $t$, and $s^{\prime}\leq s$.
Note 4.5.
In the inductive step, when the restricted formula has type 2 but it cannot be determined whether it is type $2o$ or $2d$, we observe the inductive hypothesis that $\Phi_{2o}(s,t)\leq\Phi_{2d}(s,t)$ and use the bound corresponding to type $2d$.
Algorithmic aspect. It is easy to observe that our proof in fact gives a branching algorithm for enumerating the corresponding transversals. In each case, the proof identifies an anchor (a sub-formula), sets the core variables, and solves the restricted problems recursively. We will explicitly describe the algorithm in the full version of the paper.
5 Conclusions
Our results suggest several obvious open problems. We have only managed to prove optimal bounds for monotone $\Sigma_{3}^{3}$ for $t$-threshold functions for $t\leq n/2$. It would be desirable to extend these results for general $\Sigma_{3}^{3}$. It is perhaps more challenging to extend the current techniques for $\Sigma_{3}^{k}$ for $k\geq 4$ to settle the question regarding the complexity of Majority function. There are several obstacles. It is not clear whether the modularity principle holds for the Boolean threshold functions for $k$-CNF with $k\geq 4$. Even if it holds, it is not clear how to uniformly (with respect to $k$) find the building blocks.
It would also be interesting to extend the construction for thresholds beyond $n/2$ even for monotone $\Sigma_{3}^{3}$. While we have certain guesses for the extremal constructions, proving their optimality may involve the resolution of famous hypergraph Turán problems.
It is also interesting to find a shorter version (or a simpler alternative) of our proof. It is conceivable an intelligent computer search might be able to discover a shorter version of the current proof.
Acknowledgments: We want to thank Pavel Pudlák for helpful discussions.
6 Proof of theorem˜4.1
In the rest of the paper, we present the proof of Theorem 4.1.
6.1 Configurations
To prove upper bounds on the number of $t$-transversals of a $t$-threshold monotone $k$-CNF, our arguments will be based on whether certain configurations of clauses are present in the CNF. In the following, we catalog these configurations for use in our proofs.
Definition 6.1 (Triangle configuration).
We say that a monotone CNF contains a triangle configuration if it contains three clauses of the form $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ where the clause $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$.
Definition 6.2 ( $\mathbf{\exists}$} configuration).
We say that a CNF contains a $\mathbf{\exists}$} configuration if it contains four clauses of the form $\{a,b,c\}$, $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ where the set $\{d,e,f,g,h,i\}$ is disjoint from the set $\{a,b,c\}$. We call $\{a,b,c\}$ the configuration’s spine and $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ its arms.
We sometimes regard a collection of monotone $2$-clauses as a graph where each $2$-clause is an edge and use graph-theoretic terminology to describe configurations.
Definition 6.3 (Path configuration).
We say that a collections of monotone $2$-clauses is a path of length $\ell$ if it contains $\ell$ clauses of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{\ell}\}$ where $a_{0},a_{1},\ldots,a_{\ell}$ are distinct variables.
Definition 6.4 (Cycle configuration).
We say that a collection of $2$-clauses is a cycle of length $\ell$ if it contains $\ell$ clauses of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{0}\}$ where $a_{0},a_{1},\ldots,a_{\ell-1}$ are distinct variables.
@@ 6.2 Proof of the upper bounds for $\Phi_{0}(s,t)$
There exists a pair of variables that appears in at least three distinct $3$-clauses.
@@
There exists a variable that appears in at least three distinct $3$-clauses.
@@
There exists a pair of variables that appears in two distinct $3$-clauses.
@@
There exists a variable that appears only in one clause.
@@
If a monotone $3$-CNF $\mathbb{F}$ does not satisfy the conditions $\mathscr{P}(0,2),\mathscr{P}(0,3)$ and $\mathscr{P}(0,4)$, then the following property holds: Let $\{a,b,c\}$ be any clause in the system. For $x,y\in\{a,b,c\}$ and $x\neq y$, if there exists clausess $S_{x}$, $S_{y}$ where $S_{x}$ contains $x$ and $S_{y}$ contains $y$ and $S_{x}$ and $S_{y}$ are distinct from $\{a,b,c\}$, then $S_{x}$ and $S_{y}$ are disjoint.
@@
If a monotone $3$-CNF $\mathbb{F}$ does not satisfy the condition $\mathscr{P}(0,2)$, then every variable appears at most twice. If, in addition, the property $\mathscr{P}(0,3)$ does not hold, then by lemma˜6.5, either there exists an $\mathbf{\exists}$} configuration or at least one variable appear in exactly one clause of $\mathbb{F}$. ∎
Note 6.7.
$\mathscr{P}(0,1)$ is redundant in the sense that every $\mathbb{F}\in{\Xi}_{0}$ satisfies one of the properties, $\mathscr{P}(0,i)$ for $2\leq i\leq 6$. However, we prioritize the application of induction based on $\mathscr{P}(0,1)$ to avoid dealing with restricted CNFs of three $2$-clauses with a common variable.
6.2.1 $\mathscr{P}(0,1)$ – There exists a pair of variables that appears in at least three distinct $3$-clauses
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,1)$. Let the distinct variables $a$ and $b$ appear together in the clauses $\{a,b,c\},\{a,b,d\}$ and $\{a,b,e\}$ where $c$, $d$, and $e$ are distinct variables. Either one of $\{a,b\}$ is such that it always appears in a clause together with the other variable in $\{a,b\}$ or not. We consider each case in turn and work with the corresponding anchors.
@@
Core variables Forced variables Restricted CNF parameters Inductive bound
@@ 11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{18}$ $\frac{3}{7}$
6.2.1.2 $a$ and $b$ each appear in a clause without the presence of the other variable.
We now extend the anchor by including clauses containing either $a$ or $b$ without the other variable being present. Observe that when $a=0$ and $b=1$ in the following table, we guaranteed the existence of at least one $2$-clause in the resulting restricted CNF.
Core variables Forced variables Restricted CNF parameters Inductive bound
@@
6.2.2 $\mathscr{P}(0,2)$ – There exists a variable that appears in at least three distinct $3$-clauses
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,2)$. Let $a$ be an variable that appears in three clauses $\{a,b,c\}$, $\{a,d,e\}$ and $\{a,f,g\}$ of $\mathbb{F}$ where $\{b,c\}$, $\{d,e\}$ and $\{f,g\}$ are distinct and do not have a common variable.
Core variables Forced variables Restricted CNF parameters Inductive bound
@@ 6.2.2 $\mathscr{P}(0,2)$ – There exists a variable that appears in at least three distinct $3$-clauses
6.2.3 $\mathscr{P}(0,3)$ – There exists a pair of variables that appears in two distinct $3$-clauses
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,3)$. Assume that $\mathbb{F}$ has a pair $\{a,b\}$ of variables which appears in exactly two clauses $\{a,b,c\}$ and $\{a,b,d\}$ since otherwise if $\{a,b\}$ were to appear in three clauses, we would have applied induction using $\mathscr{P}(0,1)$. We also know that $a$ (and $b$) cannot appear in any other clause since otherwise we would have applied induction using $\mathscr{P}(0,2)$.
Core variables Forced variables Restricted CNF parameters Inductive bound
@@ 6.2.4 $\mathscr{P}(0,4)$ – There exists a triangle configuration
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,4)$. Let $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ be a triangle configuration in $\mathbb{F}$ where the clause $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$. Moreover, we assume that $a$, $b$, and $c$ do not appear in any other clause of $\mathbb{F}$. It is not the case that $d=e=f$ since $\mathscr{P}(0,3)$ is not applicable. However, two of the three variables $\{d,e,f\}$ can be the same. For the last three restrictions in the following table, $n^{\prime}$ could be n5n-5. However, due to monotonicity with respect to $n$, we can replace n5n-5 with n4n-4 without loss of generality.
Core variables Forced variables Restricted CNF parameters Inductive bound
@@ 6.2.5 $\mathscr{P}(0,5)$ – There exists an $\mathbf{\exists}$} configuration
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,5)$. Let $\{a,b,c\}$ be the spine of the $\mathbf{\exists}$} configuration, and $\{a,d,e\},\{b,f,g\}$ and $\{c,h,i\}$ are its arms where the set $\{d,e,f,g,h,i\}$ is disjoint from $\{a,b,c\}$. We assume $a,b$, and $c$ do not appear in any other clauses.
Core variables Forced variables Restricted CNF parameters Inductive bound
@@
6.2.6 $\mathscr{P}(0,6)$ – There exists a variable that appears only in one clause
Assume $\mathbb{F}\in{\Xi}_{0}(s,t)$ satisfies $\mathscr{P}(0,6)$. Let $a$ be such that $a$ appears exactly once in $\mathbb{F}$ in the clause $\{a,b,c\}$. We consider two cases: Each of $b$ and $c$ appear in a different clause or one of them does not appear in any other clause.
6.2.6.1 Each of $b$ and $c$ appear in a different clause:
@@
Core variables Forced variables Restricted CNF parameters Inductive bound
@@
6.2.6.2 One of $\{b,c\}$ does not appear in any other clause:
Assume that $\mathbb{F}$ contains $\{a,b,c\}$ and $a$ and $b$ do not appear anywhere else. Either $c$ does not appear in any other clause or there is a clause of the form $\{c,d,e\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
@@ 11 $a=b=0$ n3n-3 t1t-1 $s$ 0 $\frac{1}{3}$ $\frac{1}{3}$
When $c=0$, the number of transversals when $c$ appears in another clause is dominated by the number of transversals when $c$ does not appear in any other clause.
@@ 6.3 Proof of the Upper bounds for $\Phi_{1}(s,t)$
We cover $\mathbb{F}\in{\Xi}_{1}(s,t)$ with the following properties.
@@
At least one of the variables in $\{a,b\}$ appears at least twice.
6.3.1 $\mathscr{P}(1,1)$ – At least one variable in $\{a,b\}$ appears exactly once
Assume $\mathbb{F}\in{\Xi}_{1}(s,t)$ satisfies $\mathscr{P}(1,1)$. Without loss of generality, assume that the variable $a$ appears only in the clause $\{a,b\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 $b=0$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
0 $b=1$ n2n-2 t1t-1 s1s-1 0 $\frac{7}{15}$ $\frac{1}{2}$
@@
6.3.2 $\mathscr{P}(1,2)$ – At least one of the variables in $\{a,b\}$ appears at least twice
Assume $\mathbb{F}\in{\Xi}_{1}(s,t)$ satisfies $\mathscr{P}(1,1)$. In this case, $\mathbb{F}$ has a clause of the form $\{a,c,d\}$ where $\{c,d\}$ is disjoint from $\{a,b\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@ 6.4 Proof of the Upper bounds for $\Phi_{2o}(s,t)$
We now prove the upper bounds for monotone $3$-CNFs which contain two overlapping $2$-clauses as stated in Theorem 4.1. Let $\mathbb{F}\in{\Xi}_{2o}(s,t)$ be a monotone $3$-CNF over $n\geq 3$ variables with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains the overlapping $2$-clauses $\{a,b\}$ and $\{a,c\}$.
We will first assume $s$ is odd and show that the number of $t$-transversals is at most $(\frac{5}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
11 n1n-1 t1t-1 s2s-2 0 $\frac{7}{10}$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{3}{10}$
@@
For the rest of the section, let $s$ be even. We prove that the number of $t$-transversals of $\mathbb{F}$ is bounded by $(\frac{2}{3})(\frac{2}{3})^{\frac{s}{2}}3^{t}$. For this purpose, we cover ${\Xi}_{2o}(s,t)$ with the following properties. Observe that if at least one of $b$ or $c$ appears at least three times, then one of $\mathscr{P}(2o,1),\ldots,\mathscr{P}(2o,8)$ holds. The remaining two properties cover the case when both $b$ and $c$ appear at most twice.
@@
There are two clauses containing both $b$ and $c$.
@@
There is exactly one clause in which both $b$ and $c$ appear together and exactly two other clauses, distinct from $\{a,b\}$ and $\{a,c\}$, where one contains $b$ but not $c$ and the other $c$ but not $b$.
@@
There exist clauses $\{b,d,e\}$ and $\{c,d,e\}$ where $d$ and $e$ are distinct from $a$, $b$ and $c$ and another clause that contains $b$ but not $c$ which is distinct from $\{a,b\}$ and $\{b,d,e\}$ .
@@
For each $x\in\{b,c\}$, there are exactly two clauses other than $\{a,b\}$ and $\{a,c\}$ which contain $x$ but do not contain $\{b,c\}-\{x\}$.
@@
One of $b$ or $c$ appears exactly in three clauses which do not contain the other variable and the other variable appears exactly in two clauses which do not contain the first variable.
@@
There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and exactly one clause that contains both $b$ and $c$.
@@
There are exactly two clauses that contain $b$ and exactly two clauses that contains $c$ and $b$ and $c$ do not appear together.